# Optimizing the interaction geometry of inverse Compton scattering x-ray sources

### v1

@@ III.1 Single electron with a pencil laser beam
where we used $\gamma\gg 1$ in the approximation on the right.
Figure 7 shows how $\zeta$ changes with the interaction angle for various aspect ratios, $\sigma_{L\parallel}/c\sigma_{t}$. In a head-on geometry $(\theta_{L}=\pi)$, it is obvious that $\zeta=1$. For $\theta_{L}=0$ the laser pulse is propagating in the same direction as the electron beam and again $\zeta=1$. However, for $0<\theta_{L}<\pi$, the interaction time strongly depends on $\theta_{L}$. The effect of $\zeta$ is most extreme at small interaction angles, a grazing angle geometry, with a minimum at $\theta_{L}=\cos^{-1}(\beta)\simeq 1/\gamma$.
The effect of the reduced overlap can be overcome by either applying a pulse front tilt to the laser that corrects for the tilted front in the co-moving frame [steiniger2014optical, potylitsyn2023crab, schaap2025shallow]; or by utilizing a line focused laser pulse, where $\sigma_{L\parallel}\gg\sigma_{L\perp}$. In this work we will analyze the second solution.
As an example, a 100 fs laser pulse focused to $\sigma_{L}=2.4\ \mathrm{\mu m}$ interacting under $\theta_{L}=10^{\circ}$ at $\gamma=50$, gives rise to a reduction of the interaction time with $\zeta=0.007$. As a result the brilliance is reduced by more than 4 orders of magnitude. Utilizing a line focus, with aspect ratio $\sigma_{L\parallel}/\sigma_{L\perp}=100$, we find that $\zeta=0.58$. While the laser intensity is now lower (by a factor of 100), the x-ray brilliance is almost two orders of magnitude higher than if a tight symmetrical focus was applied.
The blue curve in Fig. 7 shows that also for small deviations from $\theta_{L}=\pi$ the interaction time can become significantly reduced when utilizing a tight laser focus. This suggest that in a head-on geometry, it may pay off to approach $\theta_{L}=\pi$ as closely as possible, for instance by using parabolic mirrors with through holes [zhang2024experimental].
Figure 7: $\zeta$ as a function of the interaction angle for three different ratios of $\sigma_{L\parallel}/c\sigma_{t}$. The solid curves correspond to $\gamma=10$, whereas the dashed curves correspond to $\gamma=50$.
In Sec. III.4 we will find expressions for the optimized laser pulse dimensions. To find these, however, we first need to describe the effect of focused laser pulse for arbitrary interaction angles.
@@ III.2 Single electron with a focused laser beam
Similar to Sec. II.2 we have found that a tight laser focus is required to optimize the ICS interaction. Therefore we need to consider the effect of a focused laser pulse. In contrast to Sec. II.2, we now only have a tight focus in the y-direction. Additionally, since the interaction angle is no longer parallel to the electron velocity, the electron no longer traverses the complete longitudinal profile of the laser pulse. The effect of a focused laser pulse for grazing angle ICS is analyzed using the substitution $A(\tau)\rightarrow A(\tau)/\sqrt{1+\tau^{2}/\tau_{R}^{2}}$, similar to the method presented in [schaap2023ponderomotive]. Here we have introduced a proper Rayleigh time,
@@
$\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}=\frac{\beta\cos\theta_{L}}{1-\beta\cos{\theta_{L}}}(\sigma_{\theta L\parallel}^{2}+\sigma_{\theta L\perp}^{2}).$ (28)
@@
$\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}=1/\left(\omega_{L}\sigma_{t}\zeta\psi\right).$ (29)
@@
$\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\omega_{L}}=1/\left[2\omega_{L}\sigma_{t}\zeta\psi\ \mathcal{K}_{0}(\psi^{2}/4)\right].$ (30)
@@
Other bandwidth contributions can be included in the same method as discussed Sec. II.2. For the laser parameters under consideration, the bandwidth due to nonlinear broadening can be neglected due to the use of the line focus.
@@ III.3 Focused electron and laser beams
Besides the well known bandwidth contribution of the electron angular spread due to the spread in emission angle, $\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta\mathrm{x}}}=\gamma^{2}\sigma_{\theta e}^{2}$, in a grazing angle geometry we also have to take into account that the electron angular spread changes the interaction angle. This creates an x-ray bandwidth contribution
@@
The approximation on the right hold for $\theta_{L}\ll 1$.
The last effect we have to take into account is interaction length, $\gamma\beta c\tau_{\mathrm{int}}$, becoming longer than to the electron beta function, $\hat{\beta}=\epsilon_{n}/(\gamma\sigma_{\theta e}^{2})$. In this regime the x-ray source size varies during the interaction and we can no longer use the electron beam size in the focus as the x-ray source size. Instead we have to use the effective x-ray source size, obtained by a laser intensity- weighted average of the electron beam size:
@@ III.4 Optimization of a grazing angle interaction geometry
We now have all ingredients in place to optimize the interaction geometry of grazing angle ICS. To arrive at closed form analytic expressions, we first consider the interaction of a single electron with a focused laser pulse and optimize the peak normalized spectral angular density. Next we add the contributions for a realistic electron beam to calculate the corresponding x-ray brilliance. As a final step, we investigate the behavior for small interaction angles $\theta_{L}\ll 1$, to find the optimized electron energy. This method yields analytical expressions for the laser pulse dimensions, interaction angles and electron energy, for a desired x-ray energy.
For the interaction of a single electron, we optimize the peak normalized spectral angular density given by Eq. (31), to find optimized values of $\sigma_{L\parallel}$ and $\sigma_{L\perp}$. The dependency on the laser pulse dimensions is implicit in $A_{0}$ and $\zeta$, given by Eqs. 21 and 23 respectively. We now consider a line focused geometry such that $\psi$ is given by Eq. (27), and recognize that $A_{0}^{2}\propto 1/\sigma_{L\parallel}\sigma_{L\perp}$. We now find that function that the peak normalized spectral angular density is scales as
@@
Substitution of $\zeta=\sqrt{1/3}$ and $\psi=\psi_{\mathrm{max}}$ into Eq. (27), we also find an expression for the optimized $\sigma_{L\perp}$:
@@
The red curves in Fig. 8 show the optimized laser spot sizes $\sigma_{L\parallel}$ and $\sigma_{L\perp}$ for the generation of 500 eV x-rays under different interaction angles, according to Eqs. 39 and 40. The figure shows that for small interaction angles the optimal laser spot sizes become large in both direction, however the line focus is still maintained. The dashed green curve shows $\sigma_{e}$ for $\epsilon_{n}=$ 200 nmrad and $\sigma_{\theta}$ = 1 mrad. Since the optimized laser spot sizes can become very small outside of the grazing angle geometry, $\sigma_{e}$ should always be used as a lower bound to $\sigma_{L}$.
Figure 8: Optimized laser spot sizes (red), $\sigma_{L\parallel}^{\mathrm{opt}}$ and $\sigma_{L\perp}^{\mathrm{opt}}$, for the generation of 500 eV x-rays at varying interaction angles. The green curve shows the electron spot size at this interaction angle.
@@
$B_{\mathrm{x}}=\frac{fN_{e}}{\pi\sigma_{e}^{2}(1+\kappa^{2}/2)}\frac{1}{1+\Sigma}\omega_{\mathrm{x}}\mathcal{S}(\omega_{\mathrm{x}},0).$ (41)
Here, we have introduced $\kappa~=~\gamma^{2}\sigma_{\theta e}^{2}c\tau_{\mathrm{int}}/\epsilon_{n}$ which corrects for the effective spot size over the interaction length. For arbitrary interaction angles and beam dimensions the parameters $\zeta,\psi$ and $\Sigma$ should be calculated using Eqs. 23, 25 and 33, respectively. However, in the optimized grazing angle geometry, $\zeta=\sqrt{1/3}$ , $\psi=1.91$, and $\Sigma$ is becomes
@@
Figure 9: X-ray brilliance for the generation of 500 eV x-rays as the interaction angle is varied and the laser spot sizes are $\sigma_{L\parallel}^{\mathrm{opt}}$ and $\sigma_{L\perp}^{\mathrm{opt}}$. The inset shows a zoom in between 0-30 degrees, where the arrow denotes interaction angle corresponding to the optimized electron energy $\gamma_{\mathrm{opt}}$
Figure 9 shows x-ray brilliance for the generation of 500 eV x-rays as the interaction angle is varied and the laser pulse dimensions are changed along according to Eqs. (39) and (40). The electron beam $\epsilon_{n}=$ 200 nmrad and $\sigma_{\theta}$ = 1 mrad. The electron energy is determined by the interaction angle and the fixed x-ray energy of 500 eV. Here $\sigma_{e}$ is used as a lower bound for $\sigma_{L\parallel}$ and $\sigma_{L\perp}$.
The figure clearly shows how the brilliance increases for smaller interaction angles. However, because the effective source size becomes large at very small interaction angles, the brilliance drops significantly in this range. This results in an optimal interaction angle of equivalent interaction angle. To find an expression this optimized electron energy, we analyze how $\sigma_{L\parallel}^{\mathrm{opt}}$ and $\sigma_{L\perp}^{\mathrm{opt}}$ scale for large electron energy. By series expansion of Eq. (1) for $\theta_{L}\ll 1$ we find that
@@
Here we also find the aspect ratio of the line focus to be $\sigma_{L\parallel}/\sigma_{L\perp}\simeq 2.15\sqrt{\omega_{L}\sigma_{t}/\psi_{\mathrm{max}}}$. Since the laser spot sizes in both directions become linear with the electron energy, the peak normalized spectral angular density becomes constant w.r.t. $\gamma$. As a consequence the x-ray brilliance scales according to
@@
Using Eq. (46) with (1) the corresponding interaction angle is found. As an example for the parameters used in Fig. 9, the optimized electron energy $\gamma=103$ and the corresponding optimized interaction angle $\theta_{L}=11.3^{\circ}$. This agrees well to the result shown in Fig. 9.
Finally Eqs. (39) and (40) are used to find the optimized laser pulse dimensions, resulting the optimized grazing angle interaction geometry.
@@ III.5 Performance of a grazing angle soft x-ray Compton source
As an example, we apply the developed framework for the design of a soft x-ray Compton source with continuous tunability between 0.25 keV and 2.0 keV. This energy range has strong interest in life sciences [kordel2020laboratory, weinhardt2025soft] and catalysis [cao2020emerging, beaumont2020soft]. In this energy range achromatic reflective x-ray optics with focal lengths below 1 meter are available [kimura2022soft], which could keep the footprint within the scale of a university lab.
@@
Figure 10: X-ray brilliance for varying x-ray energies for three different geometries: head-on (red) , grazing angle with full optimization (blue), and grazing angle with constant electron energy (green). a) shows the results for soft x-rays (0.25-2 keV) and b) shows the same result with the extension up to 20 keV.
@@
The electron beamline is based on a high repetition rate C-band photoinjector [lucas2023toward]. This injector creates $Q=200$ pC electron bunches with normalized emittance of $\epsilon_{n}=200$ nmrad $(\eta=14\mathrm{\ nmrad/\sqrt{pC}})$ and 0.2% energy spread at a repetition rate of $f=1$ kHz. Subsequent LINACs can be used to accelerate the electrons to the required energy. The electrons are focussed to the interaction point with an angular spread of $\sigma_{\theta e}=1$ mrad. The interaction laser is a 100 fs, 5mJ, Yb-fiber system with a central wavelength at 1030 nm, also operating at 1 kHz [LightConversion_Pharos]. We assume the electron pulse length to be matched to the laser pulse length.
@@
We compare the x-ray source performance under three geometries. First, a full optimization as described in the previous section, where the electron energy, the interaction angle, and laser pulse shape are optimized. Second, an optimization of only the interaction angle, and laser pulse shape. Here the electron energy is constant. This has the advantage of no re-alignment of the electron optics, at the slight cost of performance. Finally, a head-on geometry optimized according to Sec. II. The corresponding x-ray brilliance throughout the x-ray energy range in the different geometries is shown in Fig. 10a. The dots show a comparison with particle tracking simulations (General Particle Tracer [PulsarPhysicsGeneralGPT]) and subsequent numerical calculation of the Liénard-Wiechert potentials [thomas2010algorithm].
@@
We find here that the grazing angle geometry is able to improve the brilliance of an ICS source by almost an order of magnitude in the soft x-ray regime. Additionally, the full optimization only differs slightly from the results of $\gamma=100$. Fig. 10b shows the performance with the x-ray energy range continued to 20 keV. We find here that the optimized head-on geometry outperforms the grazing angle geometry above 7 keV for the considered parameters.
@@
The performance of the grazing angle geometry could be improved by utilizing an appropriate pulse front tilt, such that $\zeta\rightarrow 1$. The x-ray brilliance would in this case increase by approximately the aspect ratio of the laser pulse focus $\sigma_{L\parallel}/\sigma_{L\perp}\simeq 2.15\sqrt{\omega_{L}\sigma_{t}/\psi_{\mathrm{max}}}$.
@@
Additionally the performance of the head-on geometry can be increased by using an appropriately chirped laser pulse to eliminate the effect of nonlinear broadening, which would reduce $\chi\rightarrow 0$.

### v2

@@ III.1 Single electron with a pencil laser beam
where we used $\gamma\gg 1$.
Figure 7 shows how $\zeta$ depends on the interaction angle $\theta_{L}$ for various aspect ratios $\sigma_{L\parallel}/c\sigma_{t}$. In a head-on geometry $(\theta_{L}=\pi)$, it is obvious that $\zeta=1$. For $\theta_{L}=0$ the laser pulse propagates in the same direction as the electron beam and again $\zeta=1$. However, for $0<\theta_{L}<\pi$, the interaction time strongly depends on $\theta_{L}$. The effect on $\zeta$ is most extreme at small interaction angles, a grazing angle geometry, with $\zeta$ going though a minimum at $\theta_{L}=\cos^{-1}(\beta)\simeq 1/\gamma$.
The effect of the reduced overlap can be overcome by either applying a pulse front tilt to the laser that corrects for the tilted front in the co-moving frame [39, 31, 34]; or by utilizing a line focused laser pulse, where $\sigma_{L\parallel}\gg\sigma_{L\perp}$. In this work we will analyze the second solution.
As an example, a 100 fs laser pulse focused to $\sigma_{L}=2.4\ \mathrm{\mu m}$ interacting under $\theta_{L}=10^{\circ}$ at $\gamma=50$, lowers the interaction time by a factor $\zeta=0.007$. As a result, the brilliance is reduced by more than four orders of magnitude. Utilizing a line focus, with aspect ratio $\sigma_{L\parallel}/\sigma_{L\perp}=100$, we find that $\zeta=0.58$. While the laser intensity is now lower (by a factor of 100), the x-ray brilliance is almost two orders of magnitude higher than if a tight symmetrical focus was applied.
The blue curve in Fig. 7 shows that also for small deviations from $\theta_{L}=\pi$ the interaction time can become significantly reduced when utilizing a tight laser focus. This suggest that in a head-on geometry, it may pay off to approach $\theta_{L}=\pi$ as closely as possible, for instance by using parabolic mirrors with through holes [44].
Figure 7: Reduction $\zeta$ of the overlap time as a function of the interaction angle $\theta_{L}$ for three values of the ratio $\sigma_{L\parallel}/c\sigma_{t}$ (blue, green, and yellow curves) for $\gamma=10$ (solid curves) and for $\gamma=50$ (dashed curves).
In Sec. III.4 we will find expressions for the optimized laser pulse dimensions. To find these, however, we first need to describe the effect of using a focused laser beam at arbitrary interaction angles.
@@ III.2 Single electron with a focused laser beam
Similar to Sec. II.2, we have found that a tight laser focus is required to optimize the ICS interaction. Therefore we need to consider the effect of a focused laser pulse. In contrast to Sec. II.2, we now only have a tight focus in the y-direction. Additionally, since the interaction angle is no longer parallel to the electron velocity, the electron no longer traverses the complete longitudinal profile of the laser pulse. The effect of a focused laser pulse for grazing angle ICS is analyzed using the substitution $A(\tau)\rightarrow A(\tau)/\sqrt{1+\tau^{2}/\tau_{R}^{2}}$, similar to the method presented in [36]. Here we have introduced a proper Rayleigh time,
@@
$\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}=\frac{\beta\cos\theta_{L}}{1-\beta\cos{\theta_{L}}}(\sigma_{\theta L\parallel}^{2}+\sigma_{\theta L\perp}^{2}).$ (28)
@@
$\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}=1/\left(\omega_{L}\sigma_{t}\zeta\psi\right).$ (29)
@@
$\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\omega_{L}}=1/\left[2\omega_{L}\sigma_{t}\zeta\psi\ \mathcal{K}_{0}(\psi^{2}/4)\right].$ (30)
@@
Other bandwidth contributions can be included in the same method as discussed in Sec. II.2. For the laser parameters under consideration, the bandwidth due to nonlinear broadening can be neglected due to the use of the line focus.
@@ III.3 Focused electron and laser beams
Besides the well known bandwidth contribution of the electron angular spread due to the spread in emission angle, $\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta\mathrm{x}}}=\gamma^{2}\sigma_{\theta e}^{2}$, in a grazing angle geometry we also have to take into account that the electron angular spread changes the interaction angle. This creates an x-ray bandwidth contribution
@@
The approximation on the right holds for $\theta_{L}\ll 1$.
Finally, we must take into account the effect of the interaction length, $\gamma\beta c\tau_{\mathrm{int}}$, becoming longer than the electron beta function, $\hat{\beta}=\epsilon_{n}/(\gamma\sigma_{\theta e}^{2})$. In this regime the x-ray source size varies during the interaction and we can no longer use the electron beam size in the focus as the x-ray source size. Instead we have to use the effective x-ray source size, obtained by a laser intensity-weighted average of the electron beam size:
@@ III.4 Optimization of a grazing angle interaction geometry
All ingredients which are required for optimizing the grazing angle interaction geometry are now in place. To arrive at closed form analytic expressions, we first consider the interaction of a single electron with a focused laser pulse and optimize the peak normalized spectral angular density. Next we add the contributions for a realistic electron beam to calculate the corresponding x-ray brilliance. As a final step, we investigate the behavior for small interaction angles $\theta_{L}\ll 1$, to find the optimized electron energy. This method yields analytical expressions for the laser pulse dimensions, interaction angles, and electron energy for a desired x-ray energy.
For the interaction of a single electron, we optimize the peak normalized spectral angular density given by Eq. (31), to find optimized values of $\sigma_{L\parallel}$ and $\sigma_{L\perp}$. The dependency on the laser pulse dimensions is implicit in $A_{0}$ and $\zeta$, given by Eqs. 21 and 23 respectively. We now consider a line focused geometry such that $\psi$ is given by Eq. (27), and recognize that $A_{0}^{2}\propto 1/\sigma_{L\parallel}\sigma_{L\perp}$. We now find that the peak normalized spectral angular density scales as
@@
Substituting $\zeta=\sqrt{1/3}$ and $\psi=\psi_{\mathrm{max}}$ into Eq. (27), we also find an expression for the optimized $\sigma_{L\perp}$:
@@
The red curves in Fig. 8 show the optimized laser spot sizes $\sigma_{L\parallel}$ and $\sigma_{L\perp}$ for the generation of 500 eV x-rays under different interaction angles, according to Eqs. 39 and 40. The figure shows that for small interaction angles the optimal laser spot size $\sigma_{L}$ becomes large in both directions; however, the line focus is still maintained: $\sigma_{L\parallel}\gg\sigma_{L\perp}$ over the entire angular range. The green dashed curve shows $\sigma_{e}$ for $\epsilon_{n}=$ 200 nmrad and $\sigma_{\theta}$ = 1 mrad. Since the optimized laser spot size can become very small outside of the grazing angle geometry, $\sigma_{e}$ should always be used as a lower bound for $\sigma_{L}$.
Figure 8: Optimized laser spot sizes $\sigma_{L\parallel}^{\mathrm{opt}}$and $\sigma_{L\perp}^{\mathrm{opt}}$ (red solid and dash-dotted curves) and electron spot size $\sigma_{e}$ (green dashed curve) for the generation of 500 eV x-rays as a function of interaction angle.
@@
$B_{\mathrm{x}}=\frac{fN_{e}}{2\pi\sigma_{e}^{2}(1+\kappa^{2}/2)}\frac{1}{1+\Sigma}\omega_{\mathrm{x}}\mathcal{S}(\omega_{\mathrm{x}},0).$ (41)
Here, we have introduced $\kappa~=~\gamma^{2}\sigma_{\theta e}^{2}c\tau_{\mathrm{int}}/\epsilon_{n}$ which corrects for the effective spot size over the interaction length. For arbitrary interaction angles and beam dimensions the parameters $\zeta,\psi$ and $\Sigma$ should be calculated using Eqs. 23, 25 and 33, respectively. However, in the optimized grazing angle geometry, $\zeta=\sqrt{1/3}$ , $\psi=1.91$, and $\Sigma$ becomes
@@
Figure 9: X-ray brilliance as a function of interaction angle for the generation of 500 eV x-rays, using optimized laser spot sizes $\sigma_{L\parallel}^{\mathrm{opt}}$ and $\sigma_{L\perp}^{\mathrm{opt}}$. The inset zooms in on the $0^{\circ}-30^{\circ}$ angular range, where the arrow denotes the interaction angle corresponding to the optimized electron energy $\gamma_{\mathrm{opt}}$
In Fig. 9 x-ray brilliance for the generation of 500 eV x-rays is plotted as a function of interaction angle. While the interaction angle is varied, the laser pulse dimensions are also changed, according to Eqs. (39) and (40). The normalized emittance of the electron beam is $\epsilon_{n}=$ 200 nmrad and the electron beam angular spread $\sigma_{\theta}=1$ mrad. The electron energy is determined by the interaction angle and the fixed x-ray energy of 500 eV. Here $\sigma_{e}$ is used as a lower bound for $\sigma_{L\parallel}$ and $\sigma_{L\perp}$.
The figure clearly shows how brilliance increases with decreasing interaction angles, but only up to a point. At small interaction angles, the effective source size becomes so large that the brilliance drops abruptly. This results in an optimal interaction angle, or equivalent electron energy. To find an expression for this optimized electron energy, we analyze how $\sigma_{L\parallel}^{\mathrm{opt}}$ and $\sigma_{L\perp}^{\mathrm{opt}}$ scale for large electron energy. By series expansion of Eq. (2) for $\theta_{L}\ll 1$ we find that
@@
The aspect ratio of the optimized line focus is therefore given by $\sigma_{L\parallel}/\sigma_{L\perp}\simeq 2.15\sqrt{\omega_{L}\sigma_{t}/\psi_{\mathrm{max}}}$. Since the laser spot sizes in both directions become linear with the electron energy, the peak normalized spectral angular density becomes independent of $\gamma$. As a consequence, the x-ray brilliance scales according to
@@
Combining Eq. (46) with Eq. (2) thecorresponding interaction angle is obtained. As an example for the parameters used in Fig. 9, the optimized electron energy $\gamma=103$ and the corresponding optimized interaction angle $\theta_{L}=11.3^{\circ}$. This is in good agreement with the result shown in Fig. 9.
Finally, Eqs. (39) and (40) can be used to find the optimized laser pulse dimensions, completing the full characterization of the optimized grazing angle interaction geometry.
@@ III.5 Performance of a grazing angle soft x-ray Compton source
As an example, we apply the theoretical framework described above to the design of a soft x-ray Compton source with continuous tunability between 0.25 keV and 2.0 keV. This energy range is highly relevant for life sciences [25, 43] and catalysis [11, 8]. In this energy range, achromatic reflective x-ray optics with focal lengths less than 1 m are available [24], which could keep the footprint within the scale of a university lab.
@@
Figure 10: X-ray brilliance as a function of x-ray energy for three different optimized geometries: head-on scattering (red); grazing angle scattering with optimization of both the interaction angle and the electron energy (blue); and grazing angle scattering with the electron energy kept fixed at $\gamma=100$ (green). (a) Results for soft x-rays in the range 0.25-2 keV and (b) for the entire range of 0-20 keV.
@@
The electron beamline is based on a high repetition rate C-band photoinjector [26, 2]. This injector creates $Q=200$ pC electron bunches with a normalized emittance of $\epsilon_{n}=200$ nmrad $(\eta=14\mathrm{\ nmrad/\sqrt{pC}})$ and 0.2% energy spread at a repetition rate of $f=1$ kHz. Subsequent LINACs can be used to accelerate the electrons to the required energy. The electrons are focussed to the interaction point with an angular spread of $\sigma_{\theta e}=1$ mrad. The interaction laser is a 100 fs, 5mJ, Yb-fiber system with a central wavelength $\lambda_{L}=2\pi c/\omega_{L}=1030{\,\rm nm}$, also operating at 1 kHz repetition rate [30]. We assume the electron pulse length to be equal to the laser pulse length.
@@
We compare the x-ray source performance for three geometries. First, a full optimization as described in the previous section is performed, where the electron energy, the interaction angle, and the laser pulse shape are optimized. Second, an optimization is performed in which the electron energy is kept constant and only the interaction angle and the laser pulse shape are optimized. This has the advantage that no re-alignment of the electron optics is required, while the x-ray performance is hardly affected. Finally, a head-on geometry is optimized as described in Sec. II. The resulting analytically optimized x-ray brilliances in the three different geometries are plotted with solid curves in Fig. 10a for the 0.2520.25-2 keV soft x-ray energy range and in Fig. 10b for the entire 0200-20 keV x-ray energy range. Also shown are the results of charged particle tracking simulations (General Particle Tracer [32]) combined with the numerical calculation of the Liénard-Wiechert potentials [41], which are indicated by dots. The analytical model results are in excellent agreement with the numerical calculations.
@@
We find that the grazing angle geometry is able to improve the brilliance of an ICS source in the soft x-ray regime by almost an order of magnitude, compared to the head-on geometry. In addition, we find that the optimization of the grazing angle geometry is hardly affected if the electron energy is not optimized and kept fixed at $\gamma=100$. Figure 10b shows that the optimized head-on geometry outperforms the grazing angle geometry only at x-ray energies exceeding 7 keV for the parameters considered.
@@
The performance of the grazing angle geometry could be improved even further by using an appropriate pulse front tilt such that $\zeta\rightarrow 1$. In this case, the x-ray brilliance would increase by approximately the aspect ratio of the laser pulse focus $\sigma_{L\parallel}/\sigma_{L\perp}\simeq 2.15\sqrt{\omega_{L}\sigma_{t}/\psi_{\mathrm{max}}}$.
@@
Additionally, the performance of the head-on geometry can be increased by using an appropriately chirped laser pulse to eliminate the effect of nonlinear broadening, which would reduce $\chi\rightarrow 0$.